OpenStudy (anonymous):

In a recent survery of 440 people, 1/4 said that they wached televisionevery evening, 2/5 said that they watched five or six nights each week and the remainder said they four nights or less. q1: how many people in the survery watched television every evening q2: how many people surveryed watched television five or six nights each week q3: how many people surveryed watch television four nights each week or less. please show work.

OpenStudy (anonymous):

Every evening is 1/4 of the people so 1/4 of 440 is 1/4 * 440 = 110 people five or six nights is 2/5 of the people so 2/5 of 440 is 2/5 * 440 = 176 people four nights or less - two ways to do this one. If there are 440 people then 440 - 110 - 176 = 154 people or 1/4 + 2/5 = 5/20 + 8/20 = 13/20 so the rest of the 20/20 is 7/20 so 7/20 of the people are 4 or less so 7/20 of 440 is 7/20 * 440 = 154 people